OpenStudy (ujjwal):

A parallel plate air-capacitor of capacitance C is subjected to potential difference V. What is the force between the plates of the capacitor if the separation of plates is d? a)\[\frac{CV^2}{2d}\] b)\[\frac{C^2V^2}{2d^2}\] c)\[\frac{C^2V^2}{d^2}\] d)\[\frac{C^2V}{d}\]

OpenStudy (anonymous):

q = cv. f = q.q/ r^2

OpenStudy (ujjwal):

That doesn't help. And how do you get f=q.q/r^2? I guess\[F=\frac{q\times q}{4\pi \epsilon_0 r^2}\]

OpenStudy (ujjwal):

@telltoamit ??

OpenStudy (anonymous):

oops sorry, then the answer shud be C2V2/4πϵ0d2

OpenStudy (anonymous):

anyways gtg, hope u find someone to help

OpenStudy (ujjwal):

ok.

OpenStudy (anonymous):

a) CV^2 /2d

OpenStudy (ujjwal):

How did you do it? But it says correct answer is d.

OpenStudy (ujjwal):

@Vaidehi09 Is that link reliable? well, i had also thought the same way earlier but then the answer is not correct probably..

OpenStudy (anonymous):

since the answer is not a....there must be something wrong....i'll try searching more.

OpenStudy (anonymous):

the force on the the plates is F = 1/2 QE........I just saw Prof. Lewin's lecture. so if we sub E = V/d and Q = CV then we should get F = CV^2/2d......this should be the answer.

OpenStudy (ujjwal):

yeah, may be my book is mis-printed...

OpenStudy (ujjwal):

what do you say @apoorvk ?

OpenStudy (anonymous):

here's the link for the lecture. listen to the first part. he explains why there is 1/2 in the formula. http://www.physicsforums.com/showthread.php?t=292158

OpenStudy (apoorvk):

All my latex work crashed D: (btw option A is correct, let me just tell you where you're going wrong)

OpenStudy (ujjwal):

So we are going wrong and yet we are ending with correct answer? @apoorvk

OpenStudy (anonymous):

i'll check out ur explanation tomorrow morning. its quite late now, so gtg!

OpenStudy (apoorvk):

wait option A is NOT right am sorry.

OpenStudy (ujjwal):

Are you replying @apoorvk ? I am feeling sleepy...

OpenStudy (apoorvk):

Gimme a moment please, typing it all out again can get pretty boring.

OpenStudy (ujjwal):

Well walter lewin says Force between two parallel plates is:\[F=\frac{1}{2}qE\]so, i guess, it should be (a).

OpenStudy (apoorvk):

Yeah. so... C = Q/V = ϵoA/d so, Q = CV = ϵoAV/d Now, we calculate the electric field (E) between the capacitor plates. Applying Gauss's Law to the positive plate, (let are of each face be 'A') E'.(2A) = Q/ϵo = ϵoAV/dϵo = AV/d or, E' = AV/2Ad = V/2d Now for the contribution to the net electric field bet ween the two plates would be the same in magnitude as well as direction (i.e. towards the negative plate), so the net Elctric field between the plates = E = V/2d + V/2d = V/d

OpenStudy (apoorvk):

Now. what is the force acting upon??? The question asks about the "force acting between the plates", but on which particle? The force would depend on the charge carried by the particle on which it acts!

OpenStudy (apoorvk):

Oh do you mean to ask about the attractive force between the two plates of the capacitor?

OpenStudy (ujjwal):

Its asking for the force between two plates as clearly stated in the question.

OpenStudy (experimentx):

not sure though